OpenStudy (anonymous):

simplify (-5m^2q)2(-3m^3q)3

OpenStudy (anonymous):

(-5m^2q)^2(-3m^3q)^3 (-5m^4q^3)(-3m^6q^4) -5m^4*-3m^6=15m^10 q^3*q^4=2q^7 (-5m^2q)^2(-3m^3q)^3=15m^10*2q^7

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

Welcome! God Bless! (: